# PolypDB: A Curated Multi-Center Dataset for Development of AI Algorithms in Colonoscopy

### v1

@@ IV-A Impact of multi-modality and multi-center data
One of the key strengths of PolypDB is its inclusion of data from five distinct imaging modalities—BLI, FICE, LCI, NBI, and WLI—collected from three different medical centers across Norway, Sweden, and Vietnam. This diversity is crucial in ensuring that models trained on PolypDB can generalize well across different clinical environments and patient populations. The inclusion of multi-center data helps mitigate the risk of overfitting to a specific type of imaging or patient demographic, a common challenge in medical image analysis. As our results demonstrate, models trained and evaluated on this dataset show consistent performance across different modalities, suggesting that PolypDB can serve as a valuable resource for developing more universal and robust polyp detection and segmentation models.
@@ IV-B Superior performance of PVT-CASCADE and SSFormer-L
In this study, PVT-CASCADE and SSFormer-L consistently demonstrated top-tier performance across several metrics, including mIoU, mDSC, recall, and F2 scores. Particularly in the NBI and LCI modalities, PVT-CASCADE stood out with its highest mIoU (0.7769 and 0.8344, respectively) and mDSC (0.8586 and 0.9065, respectively). This can be attributed to the powerful feature extraction capabilities of the PVTv2-B2 backbone, which effectively captures both global and local contextual information necessary for accurate polyp segmentation. The superior recall (0.9385) and F2 score (0.8941) observed in the NBI dataset further confirm PVT-CASCADE’s robustness in detecting subtle polyp structures, which is crucial in clinical settings where missing even a single polyp can have significant consequences. SSFormer-L, on the other hand, showed exceptional consistency across all modalities, particularly excelling in the WLI and LCI datasets. Its highest mIoU (0.8821) and mDSC (0.9294) in the WLI dataset, combined with its balanced performance in recall (0.9314) and precision (0.9438), underline its effectiveness in segmenting polyps with high accuracy. The integration of the MiT-PLD-B4 backbone likely contributes to this, as it enables the model to capture multi-scale information and maintain robustness across diverse visual features found in endoscopic images.
@@ IV-C Implications for clinical applications
The clinical implications of this work is substantial. By providing a publicly accessible, large-scale dataset, PolypDB enables researchers to develop more accurate and generalizable CAD systems that can assist gastroenterologists in detecting and segmenting polyps with higher precision. This can lead to a reduction in polyp miss rates, which is critical in preventing colorectal cancer. Moreover, the modality-specific benchmarks provided in this study offer guidance on selecting the most appropriate models for different imaging modalities, potentially improving the overall quality of colonoscopy procedures.
IV-D Limitation and Future directions
While PolypDB represents a significant step forward, there are limitations that warrant further exploration. For example, the dataset, while large and diverse, may still not capture the full range of variability encountered in global clinical practice. Future work could involve expanding the dataset to include more centers, imaging systems, and patient demographics. Additionally, while the current study provides a strong benchmark for segmentation and detection, there is room for improvement in addressing specific challenges, such as detecting diminutive and flat polyps, which remain difficult even for state-of-the-art models.
Despite the promising results, there are several areas for future exploration. The current models, while effective, could benefit from further refinement to enhance their robustness across all modalities. Integrating multi-modal learning, where models are trained simultaneously on multiple modalities, could improve performance by allowing the models to leverage complementary information from different imaging techniques. Additionally, incorporating real-time processing capabilities will be crucial for practical deployment in clinical settings. While this study focuses on segmentation accuracy, future work should also consider these models’ computational efficiency and speed to ensure they can be seamlessly integrated into endoscopic procedures. Finally, expanding the datasets used for training and validation, including more diverse patient populations and a wider range of polyp types, will help to ensure that these models generalize well across different clinical environments.
V Conclusion
@@
In this paper, we introduced PolypDB, a multi-center and multi-modality polyp segmentation and detection dataset designed to improve the advancement of polyp detection and segmentation in colonoscopy. PolypDB, comprising 3,934 polyp images from diverse imaging modalities and multiple medical centers, addresses the critical need for robust and generalizable data in developing CAD systems. The dataset’s diversity, in terms of both imaging modalities and geographical locations, ensures that models trained on PolypDB can perform effectively across a wide range of clinical settings, thereby enhancing their applicability in real-world scenarios. Our extensive benchmarking of SOTA segmentation and detection models demonstrated that models like SSFormer-L and YOLOv10 achieved superior performance across multiple modalities, establishing strong baselines for future research. While our study highlights the significant potential of PolypDB, it also opens up avenues for future research, including expanding the dataset’s diversity and exploring novel model architectures to address the remaining challenges in polyp detection, particularly with diminutive and flat polyps. In future research, we aim to develop a comprehensive video dataset that captures the dynamic aspects of polyp detection and segmentation during real-time colonoscopy procedures. This dataset will feature complete video sequences from multiple centers and modalities, offering richer temporal and contextual information.

### v2

@@ IV-A Impact of multi-modality and multi-center data
One of the key strengths of PolypDB is its inclusion of data from five distinct imaging modalities—BLI, FICE, LCI, NBI, and WLI—collected from three different medical centers across Norway, Sweden, and Vietnam. This diversity is crucial in ensuring that models trained on PolypDB can generalize well across different clinical environments and patient populations. The inclusion of multi-center data helps mitigate the risk of overfitting to a specific type of imaging or patient demography, a common challenge in medical image analysis. As our results demonstrate, models trained and evaluated on this dataset show consistent performance across different modalities, suggesting that PolypDB can serve as a valuable resource for developing more universal and robust polyp detection and segmentation models.
@@ IV-B Superior performance of PVT-CASCADE and SSFormer-L
In this study, PVT-CASCADE and SSFormer-L consistently demonstrated top-tier performance across several metrics, including mIoU, mDSC, recall, and F2 scores. Particularly in the NBI and LCI modalities, PVT-CASCADE stood out with its highest mIoU (0.7769 and 0.8344, respectively) and mDSC (0.8586 and 0.9065, respectively). This can be attributed to the powerful feature extraction capabilities of the PVTv2-B2 backbone, which effectively captures both global and local contextual information necessary for accurate polyp segmentation.
@@ IV-C Implications for clinical applications
PolypDB enables researchers to develop more accurate and generalizable CAD systems that can assist gastroenterologists in detecting and segmenting polyps with higher precision. This can reduce polyp miss rates, which is critical in preventing CRC. Additionally, the modality-specific benchmarks provided in this study offer guidance on selecting the most appropriate models for different imaging modalities, potentially improving the overall quality of colonoscopy procedures. The success of FL algorithm experiments demonstrated the significance of FL in colonoscopy applications, where data privacy is critical. Robust federated CAD-based algorithms (for e.g, SSFormer-L) show strong potential to enhance clinical outcomes.
V Impact of adversarial attack on PolypDB
We introduce results for the adversarial attack problem with the Fast Gradient Sign Method (FGSM) method to evaluate the robustness of deep learning models on the PolypDB dataset. By introducing small perturbations to the colonoscopy frames, we aim to expose vulnerabilities in the segmentation models. These attacks are significant in medical imaging as slight inaccuracies in the segmentation or detection results can lead to misdiagnosis. Thus, we evaluate the impact of FGSM on the baseline segmentation models on our PolypDB dataset.
Table VI shows the comparison of the qualitative results for segmentation for FSGM attack on the PolypDB dataset. From the table, we can observe that models like UNet, DeepLabV3+, DuAT, and TGANet, which performed well on the clean dataset, show substantial drops in evaluation metrics such as mIoU and DSC under adversarial conditions, demonstrating their susceptibility to perturbations. For example, UNet consistently achieved a low mIoU of 0.0485, 0.0557, 0.1057, 0.1487, and 0.0985, showing its resilience under challenging conditions.
However, models such as PVT-CASCADE and SSFormer-L achieve consistently high-performance metrics, for example, PVT-CASCADE obtained an mIoU of 0.6038 for WLI, 0.5153 for BLI, 0.4159 for FICE and 0.2560 for BLI. Similarly, SSFormer-L obtained a high mIoU of 0.5831 for WLI and 0.5476 for LCI. This underscores the resilience of these models even under challenging conditions.
Due to the high image diversity, multi-center nature and high image quality, these models still retained competitive performance. From here, we can conclude that PolypDB is not only a critical resource for benchmarking and advancing segmentation methodologies but also useful resource for developing segmentation and detection algorithms that can withstand real-world adversarial challenges such as adversarial Vulnerabilities.
TABLE VI: Comparison of quantitative results for segmentation for FGSM adversarial attack on the PolypDB dataset. The best scores are shown in bold, whereas the second best score is underlined.
Method Backbone mIoU mDSC Recall Precision F2
Dataset: PolypDB (BLI)
U-Net - 0.0485 0.0876 0.5734 0.0504 0.1634
DeepLabV3+ ResNet50 0.1698 0.2759 0.5278 0.2904 0.3584
PraNet Res2Net50 0.1159 0.1943 0.7611 0.1206 0.3254
CaraNet Res2Net101 0.2499 0.3791 0.4220 0.3701 0.3996
TGANet ResNet50 0.0760 0.1365 0.4943 0.0856 0.2204
PVT-CASCADE PVTv2-B2 0.2560 0.3424 0.4420 0.3743 0.3992
DuAT PVTv2-B2 0.2249 0.3367 0.5310 0.2677 0.4214
SSFormer-L MiT-B4 0.1962 0.2910 0.4972 0.2558 0.3677
Dataset: PolypDB (FICE)
U-Net - 0.0557 0.0996 0.4973 0.0570 0.1832
DeepLabV3+ ResNet50 0.2080 0.2942 0.4771 0.2138 0.3813
PraNet Res2Net50 0.3547 0.4320 0.5376 0.6518 0.4880
CaraNet Res2Net101 0.4061 0.4729 0.5066 0.6010 0.4904
TGANet ResNet50 0.2376 0.3249 0.5253 0.2457 0.4159
PVT-CASCADE PVTv2-B2 0.4159 0.5273 0.8009 0.4584 0.6247
DuAT PVTv2-B2 0.3263 0.4048 0.7739 0.3548 0.4805
SSFormer-L MiT-B4 0.3610 0.4670 0.7283 0.3762 0.5683
Dataset: PolypDB (LCI)
U-Net - 0.1057 0.1731 0.4809 0.1555 0.2376
DeepLabV3+ ResNet50 0.4813 0.6397 0.7441 0.6094 0.6862
PraNet Res2Net50 0.3654 0.4857 0.6395 0.5224 0.5501
CaraNet Res2Net101 0.4185 0.5500 0.6092 0.6021 0.5695
TGANet ResNet50 0.4946 0.6383 0.7308 0.6472 0.6796
PVT-CASCADE PVTv2-B2 0.4226 0.6288 0.7939 0.5434 0.7122
DuAT PVTv2-B2 0.3458 0.4691 0.7934 0.3616 0.5992
SSFormer-L MiT-B4 0.5476 0.6961 0.7934 0.7185 0.7372
Dataset: PolypDB (NBI)
U-Net - 0.1487 0.2202 0.6561 0.1785 0.3165
DeepLabV3+ ResNet50 0.1982 0.3069 0.6252 0.2867 0.4029
PraNet Res2Net50 0.3798 0.5045 0.6237 0.5989 0.5388
CaraNet Res2Net101 0.3358 0.4513 0.5584 0.5003 0.4725
TGANet ResNet50 0.2253 0.3468 0.5008 0.4154 0.3825
PVT-CASCADE PVTv2-B2 0.5153 0.6543 0.8499 0.6187 0.7198
DuAT PVTv2-B2 0.3852 0.5063 0.6483 0.4846 0.5612
SSFormer-L MiT-B4 0.2802 0.3911 0.7200 0.3242 0.4881
Dataset: PolypDB (WLI)
U-Net - 0.0985 0.1659 0.5967 0.1311 0.2563
DeepLabV3+ ResNet50 0.4930 0.6030 0.7670 0.5832 0.6585
PraNet Res2Net50 0.5732 0.6753 0.7769 0.6895 0.7067
CaraNet Res2Net101 0.5363 0.6440 0.7986 0.6327 0.6962
TGANet ResNet50 0.4413 0.5531 0.7977 0.5127 0.6293
PVT-CASCADE PVTv2-B2 0.6038 0.7070 0.8030 0.7165 0.7374
DuAT PVTv2-B2 0.5476 0.6574 0.8022 0.6553 0.7059
SSFormer-L MiT-B4 0.5831 0.6292 0.8273 0.6733 0.7386
VI Conclusion
@@
We introduced PolypDB, a multi-center and multi-modality polyp segmentation and detection dataset for advancing polyp detection and segmentation in colonoscopy. It comprises 3,934 polyp images from diverse imaging modalities and multiple medical centers and addresses the critical need for robust and generalizable data in developing CAD systems. The dataset’s diversity, in terms of imaging modalities and geographical locations, ensures that models trained on PolypDB can perform effectively across a wide range of clinical settings, thereby enhancing their applicability in real-world scenarios. Our extensive benchmarking of SOTA segmentation, detection and federated learning settings demonstrated strong baselines for future research. In upcoming work, we aim to develop a comprehensive video dataset that captures the dynamic aspects of polyp detection and segmentation during real-time colonoscopy procedures.